OpenStudy (anonymous):

A body of mass 2 kg at O has an initial velocity of 3m/s along OE and it is subjected to a force of 4N perpendicular to OE the distance covered in 4 seconds?

OpenStudy (anonymous):

@shubhamsrg @ajprincess @hartnn

OpenStudy (anonymous):

The answer should be 20

OpenStudy (shubhamsrg):

 force along OE =0 thus ,velocity in OE direction is constant = 3 force along OF =4N. mass of body=2kg. acceleration=4/2=2 distance moved by the body in OE in 4secs =vt=3.4=12m. distance moved by the body in OE in 4secs =ut+at^2/2=0+16=16m. net distance = sqrt(12^2 + 16^2) = 20